Scenario 1: Aspirin and Kidney Cancer

A group of researchers were interested in studying the use of aspiring and the incidence of kidney cancer. They designed their study and enrolled 13,987 white, affluent, and well-educated residents of a retirement community (9,621 women and 4,366 men).   3,490 men, and 6,021 women did not use aspirin upon enrollment. There were a total of 25 cases of kidney cancer that were reported after 6.5 years of follow up (15 among men, 10 among women). Among the males who developed kidney cancer, 9 used aspirin; of the females who developed kidney cancer, 4 used aspirin.

Outcome status – Kidney Cancer Present

Outcome status – Kidney Cancer Absent

Total

Exposure status – Aspirin Used (Present)

Click or tap here to enter text.

Click or tap here to enter text.

Click or tap here to enter text.

Exposure status - No Aspirin Used (Absent)

Click or tap here to enter text.

Click or tap here to enter text.

Click or tap here to enter text.

Total

Click or tap here to enter text.

Click or tap here to enter text.

Click or tap here to enter text.
Answer #1

Scenario 1: Aspirin and Kidney Cancer

Total no.of subjects = 9621 + 4366 = 13987.

Total no.of subjects did not used aspirin = 3490 + 6021 = 9511

Total no.of subjects used Aspirin = 13987 - 9511 = 4476

Total no.of subjects who developed kidney disease = 25.

Kidney cancer that developed in aspirin used subjects = 9 + 4 = 13.

Kidney cancer developed in non-aspirin used subjects = 25 - 13 = 12

Total no.of subjects not developed kidney cancer in aspirin used subjects = 4476 - 13 = 4463

Total no.of subjects not developed kidney disease in non-aspirin used subjects = 9511 - 12 = 9499

Outcome status – Kidney Cancer Present

Outcome status – Kidney Cancer Absent

Total

Exposure status – Aspirin Used (Present)

13

4463

4476

Exposure status - No Aspirin Used (Absent)

12

9499

9511

Total

25

13962

13987